# Polynomial-like dynamics of analytic maps

### v1

@@ 2. Preliminary statements
The set $\Delta:=Y\setminus Int(g^{-1}(Y))$ is a compact subset of $\Omega^{*}$, therefore, there exist $r>0$ and $\kappa\in(0,1)$ as follows: for every $a\in\Delta$, $B(a,2r)$ is a topological disk and every branch $g^{-1}:B(a,2r)\to\Omega_{*}$ contracts the distance in $\Omega^{*}$ by a factor at leat $\kappa$ (here and below in the proof $B(x,t)$ is the disk centered at $x\in\Omega^{*}$ and of radius $t>0$ in the hyperbolic metric $\rho_{*}$ of $\Omega^{*}$).
Note that $\partial Y\subset\partial\Omega_{j}\subset\Omega^{*}$ and $\partial Y^{*}$ is the disjoint union of $\partial Y$ and $\partial K$. Let $Y^{*}_{r}=Y^{*}\cup_{a\in\partial Y}\bar{B}(a,r)$.
@@
Let $\Gamma_{*}$ be the outer boundary of $Y^{*}_{r}$ (i.e, the boundary of the topological hull of $Y^{*}_{r}$). Note that $\Gamma_{*}$ is a compact subset of $\Omega^{*}$. Let $V_{*}$ be a (simply connected) domain bounded by $\Gamma_{*}$. We claim that $g^{-1}(V_{*})$ is compactly contained in $V_{*}$. It’s enough to show that $g^{-1}(\Gamma_{*})\subset V_{*}$. To this end, let $z\in\Gamma_{*}$ and $z^{\prime}\in g^{-1}(z)$. There exist $a\in\partial Y$, such that $\rho_{*}(z,a)\leq r$, and a branch $\psi$ of $g^{-1}$ in a neighborhood of $\bar{B}(a,r)$ such that $\psi(z)=z^{\prime}$. If $a^{\prime}:=\psi(a)$, then $\rho_{*}(a^{\prime},z^{\prime})\leq k\rho_{*}(a,z)\leq kr<r$. As $a^{\prime}\in g^{-1}(\partial Y)\subset\Delta$, it follows that $z^{\prime}\in Y^{*}_{r}$. Thus $V^{\prime}_{*}:=g^{-1}(V_{*})$ is compactly in $V_{*}$, i.e., $g:V_{*}^{\prime}\to V*$ is the required PL map.
(C). If $\overline{V}_{0}\subset V_{-n}$ for some $n>0$, then $V_{0}\subset g^{-n}(V_{0})\subset g^{-n}(U)\subset U^{\prime}$. Therefore, if (C) holds for every $V_{0}$, then $U\subset U^{\prime}\subset U$, i.e., $U=U^{\prime}$, a contradiction. This proves the last claim, i.e., if $\Gamma_{0}$ is close enough to $\partial U$, then (C) don’t happen. For the first claim, notice that $\overline{g^{n}(V_{0})}\subset V_{0}$. Hence, by the Wolff-Denjoy theorem (eg [1], the sequence of domains $g^{kn}(V_{0})$, $k=1,2,$..., shrink to an attracting fixed point $a\in V_{0}$ of $g^{n}$. Then $g^{i}(\overline{V_{0}})$ tend to the cycle $O\ni a$ such that $O\subset P_{g}\subset V_{0}$. Therefore, $O=g^{i}(O)\subset g^{i}(V_{0})$ for all $i\geq 0$. As $g^{i}(V_{0})\to O$, it follows that $O=\{a\}$.
We have the following
Corollary 2.3.
Assume that $g:U^{\prime}\to U$ has no a PL restriction to a neighborhood of its postcritical set $P_{g}$. Let $R_{g}$ be the closure (in the Riemann sphere $\hat{\mathbb{C}}$) of the set of repelling periodic points of $g:U^{\prime}\to U$. Then $R_{g}=\partial K_{*}$ in the case (2) while in cases (1) and (3) the compact $R_{g}\subset\hat{\mathbb{C}}$ contains a (non trivial) continuum which has common points with the range $U$ as well as with its boundary $\partial U$.
Proof.
[…]
Repeating an argument as in the case (A) of the proof above, we obtain
Lemma 2.4.
Let $a$ be a repelling periodic point of $g$ in the unbounded component $\Omega$ of $\mathbb{C}\setminus P_{g}$. If $z\in U\cap\Omega$, then for every simply connected domain $W$ in $\Omega$ that contains points $a$ and $z$ there exists a sequence of inverse branches $g^{-n_{k}}:W\to\mathbb{C}$ such that $g^{-n_{k}}(W)\to a$ uniformly on compacts.
Next claim describes properties of completely invariant sets inside of a forward invariant one.
Proposition 2.5.
Let $X$ be a compact forward invariant set such that $P_{g}\subset X$. Let
$D_{\infty}=U\setminus\hat{X}$,
where $\hat{X}$ is the topological hull of $X$. Then:
  1. (1)
    $g^{-1}(D_{\infty})\subset D_{\infty}$,
  2. (2)
    $g(\hat{X})\subset\hat{X}$.
If
$X_{*}=\{x\in\hat{X}:GO(x)\subset\hat{X}\}$
then
  1. (3)
    $X_{*}$ is the maximal completely invariant compact (perhaps, empty) which is contained in $\hat{X}$. Furthermore, $X_{*}$ is a full compact.
Finally, assuming $X_{*}\neq\emptyset$, either $P_{g}\subset X_{*}$ or $\partial X_{*}\subset P_{g}$.
Proof.
We split it into several parts.
Step I. We show that:
(1) $g^{-1}(D_{\infty})\subset D_{\infty}$.
(2) $g(\hat{X})\subset\hat{X}$.
Proof of (1). Since $X$ is forward invariant, then $X\subset\cap_{i\geq 0}U_{-i}$. It is enough to prove that, for every $i\geq 0$, $g_{i}^{-1}(U_{-i}\setminus\hat{X})\subset U_{-i-1}\setminus\hat{X}$. As $U_{-i}\setminus\hat{X}$ is a connected open set, one can join any $z\in U_{-i}\setminus\hat{X}$ with $\partial U_{-i}$ by a simple curve $\gamma$ in $U_{-i}\setminus\hat{X}$. Let $\gamma^{\prime}$ be a component of $g_{i}^{-1}(\gamma)$. It is also a simple curve that joins a preimage of $z$ to $\partial U_{-i-1}$. If $\gamma^{\prime}$ crossed $\partial\hat{X}$ then $\gamma$ would contain a point of $\hat{X}$ because $g(\partial\hat{X})\subset g(X)\subset X\subset\hat{X}$, a contradiction. Therefore, $\gamma^{\prime}\subset U_{-i-1}\setminus\hat{X}$.
(2) follows from (1) as $\hat{X}=U^{\prime}\setminus D_{\infty}$.
Define
(2.1) $X_{esc}=\{x\in\hat{X}:g^{n}(y)=x\mbox{ for some }n\in\mathbb{N},\ y\in D_{\infty}\}.$
As $D_{\infty}$ is open and $g$ is continuous, $X_{esc}$ is open (rel. $\hat{X}$): for every $x\in X_{esc}$ there exist a disk $B_{x}$ centered at $x$ and some $n(x)>0$ such that $B^{\prime}_{x}\subset D_{\infty}$, for a component $B^{\prime}_{x}$ of $g^{-n(x)}(B_{x})$. Hence, we can define the following open subset of $U$ by either of two equalities:
$D_{*}=D_{\infty}\cup\cup_{x\in X_{esc}}\cup_{n\geq 0}g^{-n}(B_{x})=D_{\infty}\cup\cup_{x\in X_{esc}}BO(x).$
Let
$X_{*}=U\setminus D_{*},$
a closed subset of $\hat{X}$. As $g(\hat{X})\subset\hat{X}$ so that $GO(x)$ is well-defined for all $x\in\hat{X}$, we have equivalently:
(2.2) $X_{*}=\{x\in\hat{X}:GO(x)\subset\hat{X}\}$.
Step II. $g^{-1}(D_{*})\subset D_{*}$ and $g^{-1}(X_{*})=X_{*}$.
Indeed, as $g^{-1}(D_{\infty})\subset D_{\infty}$, the first inclusion follows from the definition of $D_{*}$. That $X_{*}$ is completely invariant follows from (2.2).
Step III. $X_{*}$ is a full compact, i.e., $X_{*}=\hat{X}_{*}$.
[…]
(2.3) $g^{-1}(\partial\hat{X}_{*})\subset\partial\hat{X}_{*}$.
[…]
As $D_{\infty}\cap\hat{X}$$\emptyset$ while $\hat{X}_{*}\subset\hat{X}$, (2.4) implies, by the definition of $X_{*}$, that $Int(\hat{X}_{*})\subset X_{*}$. This ends the proof of Step III and claims (1)-(3) of the proposition.
Step IV. It remains to show that either $P_{g}\subset X_{*}$ or $\partial X_{*}\subset P_{g}$. So assume that $X_{*}\neq\emptyset$ and $P_{g}\nsubseteq\partial X_{*}$.
First, we claim that for every $j\geq 0$ there exists a repelling periodic point inside of $U_{-j}$ in any neighborhood of $\partial U_{-j}$. Indeed, if this is not the case for some $j$, then the alternative (B) of Proposition 2.2 must hold. As It follows now from Proposition 2.2 that since $\hat{X}\subset\cap_{i\geq 0}U_{-i}$, then $\hat{X}$ is a subset of the filled Julia set $K_{g}$ of a polynomial-like restriction of the map $g:U^{\prime}\to U$. On the other hand, as $g$ is conjugate near $K_{g}$ to a polynomial of degree $d$ which is not a power map $z^{d}$, the only completely invariant subset of $K_{g}$ is $K_{g}$ itself and then $P_{g}\subset K_{g}=X_{*}$, a contradiction.
Now we assume by a contradiction that there exists $z\in\partial X_{*}\setminus P_{g}$. Consider two cases:
(a) $z\in\partial\hat{X}$. As $\partial\hat{X}=\partial D_{\infty}$, there exists a ball $B$ centered at $z$ such that $B\subset U_{-j}$ for some $j\geq 0$ and some $x\in B\cap D_{\infty}$. In particular, the family $\mathcal{G}$ of all branches of all $g^{-n}$ are well-defined univalent functions in $B$ and don’t take values in the set $\cup_{k\geq 0}g^{-k}(P_{g})$. Hence, $\mathcal{G}$ is a normal family. On the other hand, we proved that there exists a repelling periodic point $a_{j}\in D_{\infty}\cap U_{-j}$. As $x\in B\cap D_{\infty}$, by Lemma 2.4, there exists a sequence of branches $g^{-n_{k}}\in\mathcal{G}$ such that $g^{-n_{k}}\to a_{j}$ uniformly on a compact in $B$ that contains $z\in X_{*}$. But this is a contradiction as $X_{*}$ is completely invariant and away from $a_{j}$.
[…]
@@
Assume the contrary and let $g^{p}(a)=a$ and $|\lambda|>1$ for some $p\in\mathbb{N}$, $a\in U^{\prime}$ where $\lambda=(g^{p})^{\prime}(a)$. Replacing the map $g:U^{\prime}\to U$ by $g^{p}:g^{-p}(U)\to U$ and $d$ by $d^{p}$, we can assume that $p=1$, that is, $a$ is a repelling fixed point of $g$.
Let $\psi$ be a local branch of $g^{-1}$ that fixes $a$ and is defined in a small neighborhood $Z$ of $a$ such that $\overline{\psi(Z)}\subset Z$. For each $n>0$, $\psi^{n}(Z)\subset Z$, $\psi^{n}(a)=a$ and $(\psi^{n})^{\prime}(a)=\lambda^{-n}$. One can also assume that the Konigs linearization function $S:Z\to\mathbb{C}$ of $\psi$ maps $Z$ onto a small disk centered at zero, where $S:Z\to\mathbb{C}$ is a univalent map such that $S^{\prime}(a)\neq 0$ and $S\circ\psi=\lambda^{-1}S$ on $Z$.
@@ 3.1. The case of connected $K$
Let $\pi:\mathbb{D}^{*}:=\{|w|>1\}\to\mathbb{C}\setminus K$ be the conformal isomorphism such that $\pi(\infty)=\infty$. Let $\hat{g}:=\pi^{-1}\circ g\circ\pi:\hat{U}^{\prime}\to\hat{U}$ where $\hat{U}^{\prime}=\pi^{-1}(U^{\prime}\setminus K)$, $\hat{U}=\pi^{-1}(U\setminus K)$.
Let us fix a simply connected domain $W$ in $\mathbb{D}^{*}$ such that, if $\hat{Z}=\pi^{-1}(Z)$, then $\hat{a}\in\hat{Z}\subset W$ and $\partial W\cap S^{1}$ contains a nontrivial arc $l$ of $S^{1}$ Since all branches $\hat{g}^{-n}$ are well-defined in $W$, for every $n$, the local map $\hat{\psi}^{n}:\hat{Z}\to\hat{Z}$ admits an analytic continuation along every curve in $W$ with the initial point $\hat{a}$. Therefore, every $\hat{\psi}^{n}:\hat{Z}\to\hat{Z}$. extends to a univalent map $\hat{\psi}_{n}:W\to\mathbb{C}$. Let us "exhaust" $W$ by an increasing sequence of simply connected domains $W_{i}$, $i=1,2,$..., such that $\overline{W_{i}}\subset W$, for each $i$. As $(\hat{\psi}_{n})^{\prime}(\hat{a})=(\hat{\psi}^{n})^{\prime}(\hat{a})\to 0$, by Koebe’s distortion theorem (or simply by the normality argument), for every $i$ there exists $n_{i}$ such that $\hat{\psi}_{n_{i}}(W_{i})\subset\hat{Z}$. This allows us to extend the Konigs linearization function $\hat{S}:\hat{Z}\to\mathbb{C}$ of $\hat{\psi}:\hat{Z}\to\hat{Z}$ to a holomorphic (in fact, univalent) function $\hat{S}_{i}:W_{i}\to\mathbb{C}$ according to the equation $\hat{S}_{i}(w)=\lambda^{n_{i}}\circ\hat{S}\circ\hat{\psi}_{n_{i}}(w)$. In this way, $\hat{S}$ extends to a holomorphic function $\hat{S}:W\to\mathbb{C}$. On the other hand, the above equation implies easily that $\hat{S}(w)\to\infty$ as $w\to\partial W\cap S^{1}$. This is a contradiction with the Riesz Uniqueness theorem (or more general Privalov’s theorem), e.g. [8].
@@
We apply results of Appendices A-B. By Theorem A.1 and Corollary A.1, the logarithmic capacity of the completely invariant compact set $K$ is positive so Corollary B.1 applies. Let us fix an open subset of $\mathbb{C}$ as follows:
@@
Let $\varepsilon>0$ be the distance between the set $\cup_{i=1}^{m}\partial Q_{i}$ and $K$. For this $\varepsilon$, we find a simply connected domain $W$ as in Corollary B.1, in particular, the connected set $\pi(W)\subset\mathbb{C}\setminus K$ must lie in one and only one of the domains $Q_{1},...,Q_{m}$. W.l.o.g. one can assume that
@@
Let us joint the point $a$ and a point $q\in\partial\pi(W)\setminus K$ by a simple arc $\alpha$ which, on the one hand, lies in the range $U_{0}$ of $g$, and, on the other hand, except its end point $q$, lies outside of $\overline{\pi(W)\cup K}$ as well as outside of all others $\overline{Q_{i}}$, $2\leq i\leq m$, and also crosses the Jordan curve $\partial Q_{1}$ at a unique point. Choose a point $\hat{q}\in\partial W\cap\mathbb{D}$ such that $\pi(\hat{q})=q$ and let $\hat{\alpha}$ be the lift of the curve $\alpha$ by the covering $\pi$ that begins at $\hat{q}$. Then $\tilde{\alpha}$ is also a simple arc (in $\mathbb{D}$) with the end points $\hat{q}$ and some $\hat{a}\in\mathbb{D}$ such that $\pi(\hat{a})=a$. Note that $\hat{\alpha}\cap\overline{W}=\{\hat{q}\}$. Let us slightly "thicken" the curve $\hat{\alpha}$ to a narrow "tube" $T$ so that $\Omega:=T\cup W$ is a simply connected domain in $\mathbb{D}$ which contains the point $\hat{a}$. Then $\pi(\Omega)=\pi(T)\cup\pi(W)$ is a domain in $\mathbb{C}\setminus K$ which is disjoint with $\overline{Q_{i}}$, for all $2\leq i\leq m$. As $\alpha$ crosses $\partial Q_{1}$ at a single point, we can also assume that $Q:=\pi(\Omega)\cup Q_{1}=\pi(T)\cup Q_{1}$ is a simply connected domain, which is disjoint with all others $Q_{2},...,Q_{m}$.
Let $\psi:Z\to Z$ be the local inverse branch of $g$ introduced at the beginning of Sect. 3. As the neighborhood $Z$ of the fixed point $a$ can be chosen small enough and $\pi(\hat{a})=a$, we can "lift" the map $\psi:Z\to Z$ to a univalent map $\hat{\psi}:\hat{Z}\to\hat{Z}$ where $\hat{Z}$ is a neighborhood of $\hat{a}$, so that $\pi:\hat{Z}\to Z$ is a homeomorphism and $\pi\circ\hat{\psi}=\psi\circ\pi$ on $\hat{Z}$. For each $n\in\mathbb{N}$, consider the $n$-iterate $\hat{\psi}^{n}:\hat{Z}\to\hat{Z}$.
Claim 0. For each $n$, $\hat{\psi}^{n}:\hat{Z}\to\hat{Z}$ can be extended to a well-defined holomorphic function (which we call $\hat{\psi}_{n}$) in the domain $\Omega$.
Indeed, let $\hat{\beta}$ be any curve in $\Omega$ that starts at $\hat{a}$. Then $\beta:=\pi\circ\hat{\beta}$ is a curve in $\pi(\Omega)$ that starts at $a$ and lies in $U_{0}\setminus K$. As $g^{n}:U_{-n}\setminus K\to U_{0}\setminus K$ is an unbranched covering map, the local branch $\psi^{n}:Z\to Z$ of $g^{-n}$ has an analytic continuation along the curve $\beta$, and it is lifted by the unbranched covering map $\pi:\mathbb{D}\to\hat{\mathbb{C}}\setminus K$ to an analytic continuation of $\hat{\psi}^{n}$ along the curve $\hat{\beta}$. Since the latter is any curve starting at $\hat{a}$ in the simply connected domain $\Omega$, by the Monodromy theorem, the function $\hat{\psi}_{n}$ is well-defined in the whole domain $\Omega$ and $\hat{\psi}_{n}=\hat{\psi}^{n}$ in $\hat{Z}$.
@@
Claim 1. For each $n$, the function $\hat{\psi}_{n}:\Omega\to\mathbb{C}$ is univalent.
Indeed, assume the contrary: $\hat{\psi}_{n}(x_{1})=\hat{\psi}_{n}(x_{2})$ for some $x_{1}\neq x_{2}$ in $\Omega$. Let $\hat{l}$ be a simple curve in $\Omega$ joining $x_{1}$ and $x_{2}$. As all branches of $g^{-n}$ are inverse branches of a well-defined unbranched covering map $g^{n}:U_{-n}\setminus K\to U_{0}\setminus K$ and since $Q=\pi(\Omega)\cup Q_{1}=\pi(T)\cup Q_{1}$ is a simply connected domain, which is disjoint with some other $Q_{j}$, $j\neq 1$, the following must hold:
@@
  • an analytic continuation of some local branch of $g^{-n}$ along the closed curve $l$ maps this curve again onto some closed curve $L$,

### v2

@@ 3. On BI maps
The set $\Delta:=Y\setminus Int(g_{0}^{-1}(Y))$ is a compact subset of $\Omega^{*}$, therefore, there exist $r>0$ and $\kappa\in(0,1)$ as follows: for every $a\in\Delta$, $B_{h}(a,2r)$ is a topological disk and every branch $g_{0}^{-1}:B_{h}(a,2r)\to\Omega_{*}$ contracts the distance in $\Omega^{*}$ by a factor at leat $\kappa$ (here and below in the proof $B_{h}(x,t)$ is the disk centered at $x\in\Omega^{*}$ and of radius $t>0$ in the hyperbolic metric $\rho_{*}$ of $\Omega^{*}$).
Note that $\partial Y\subset\partial\Omega_{j}\subset\Omega^{*}$ and $\partial Y^{*}$ is the disjoint union of $\partial Y$ and $\partial K$. Let $Y^{*}_{r}=Y^{*}\cup_{a\in\partial Y}\bar{B_{h}}(a,r)$.
@@
Let $\Gamma_{*}$ be the outer boundary of $Y^{*}_{r}$ (i.e, the boundary of the topological hull of $Y^{*}_{r}$). Note that $\Gamma_{*}$ is a compact subset of $\Omega^{*}$. Let $V_{*}$ be a (simply connected) domain bounded by $\Gamma_{*}$. We claim that $g_{0}^{-1}(V_{*})$ is compactly contained in $V_{*}$. It’s enough to show that $g_{0}^{-1}(\Gamma_{*})\subset V_{*}$. To this end, let $z\in\Gamma_{*}$ and $z^{\prime}\in g_{0}^{-1}(z)$. There exist $a\in\partial Y$, such that $\rho_{*}(z,a)\leq r$, and a branch $\psi$ of $g_{0}^{-1}$ in a neighborhood of $\bar{B}_{h}(a,r)$ such that $\psi(z)=z^{\prime}$. If $a^{\prime}:=\psi(a)$, then $\rho_{*}(a^{\prime},z^{\prime})\leq k\rho_{*}(a,z)\leq kr<r$. As $a^{\prime}\in g_{0}^{-1}(\partial Y)\subset\Delta$, it follows that $z^{\prime}\in Y^{*}_{r}$. Thus $V^{\prime}_{*}:=g_{0}^{-1}(V_{*})$ is compactly in $V_{*}$, i.e., $g:V_{*}^{\prime}\to V*$ is the required PL map.
(C). If $\overline{V}_{0}\subset V_{-n}$ for some $n>0$, then $\overline{g^{n}(V_{0})}\subset V_{0}$. Hence, by the Wolff-Denjoy theorem (e.g., [2]), the sequence of domains $g^{kn}(V_{0})$, $k=1,2,$..., shrink to an attracting fixed point $a\in V_{0}$ of $g^{n}$. Then $g^{i}(\overline{V_{0}})$, $i\geq n$, tend to the cycle $O\ni a$. As $P_{g}\subset V_{0}$ is a closed invariant set, then $O\subset P_{g}\subset V_{0}$. Therefore, $O=g^{i}(O)\subset g^{i}(V_{0})$ for all $i\geq n$ where $g^{nk+j}(V_{0})$ shrinks to $g^{j}(a)$ as $k\to\infty$. Since $P_{g}\subset g^{i}(V_{)})$ for all $i\geq n$, it follows that $O=\{a\}$. If a completely invariant compact $K\subset V$ exists, then $K=\{a\}$ and $f^{-1}(a)=a$. In other words $P_{g}=\{a\}$ and $a$ is a fixed critical point of order $d$, which is excluded by Definition 1.1. Finally, if the alternative (C) holds for a sequence of such $\Gamma_{0}$’s that tends to $\partial U_{0}$, then the sequence of corresponding Jordan domains $V_{0}$’s bounded by $\Gamma_{0}$’s exhausts $U_{0}$, hence, by the above, there is an attracting fixed point $a\in\cap_{i\geq 0}g^{-i}(U_{0})$ such that $g^{n}(z)\to a$ for all $z\in U=\cup_{i\geq i}U_{0}$, i.e., $g:U^{\prime}\to U$ is trivial.
@@
Assume the contrary: $g:U^{\prime}\to U$ has a completely invariant full compact set $K$ such that $P_{g}\subset K\subset\cap_{n\geq 0}U_{-n}$, and there exist a point $a\in U_{0}\setminus K$ and sequences $n_{i}\to\infty$ and $a_{i}\to a$ such that $a_{i}\in U_{-n_{i}}$, $g^{n_{i}}(a_{i})=a_{i}$ and, for each $i$, $|\lambda_{i}|>1$ where $\lambda_{i}=(g^{n_{i}})^{\prime}(a_{i})$.
For every $i$, as $a_{i}\in U_{0}\cap U_{-n_{i}}$ and $|\lambda_{i}|>1$, there exists a branch $\psi_{n_{i}}$ of $g_{0}^{-n_{i}}$ which is defined (hence, univalent) in some neighborhood of $a_{i}$ such that $\psi(a_{i})=a_{i}$. As $a\notin P_{g}$, there exists a disk $Z:=B(a,r)\subset U_{0}$ such that $\overline{Z}\subset U_{0}\setminus P_{g}$. Therefore, each $\psi_{i}$ extends to a univalent map to $Z$. Repeating arguments as in the proof of the alternative (A), Proposition 3.2 (where one can take $V_{0}$ to be any Jordan domain such that $P_{g}\subset V_{0}\subset\overline{V_{0}}\subset U_{0}$ and $a\in\Gamma_{0}:=\partial V_{0}$) we obtain that $\lambda_{i}\to\infty$. Equivalently, $\psi_{i}^{\prime}(a_{i})=\lambda_{i}^{-1}\to 0$. In turn, shrinking the radius $r$ of the disk $Z$ if necessary, this implies that $a_{i}\in\psi_{n_{i}}(Z)\subset\overline{\psi_{n_{i}}(Z)}\subset Z$ for all $i$ big enough.
@@ 4.1. The case of connected $K$
Let $\pi:\mathbb{D}^{*}:=\{|w|>1\}\to\mathbb{C}\setminus K$ be the conformal isomorphism such that $\pi(\infty)=\infty$. Replace $g:U^{\prime}\setminus K\to U\setminus K$ by $\hat{g}:=\pi^{-1}\circ g\circ\pi:\hat{U}^{\prime}\to\hat{U}$ where $\hat{U}^{\prime}=\pi^{-1}(U^{\prime}\setminus K)$, $\hat{U}=\pi^{-1}(U\setminus K)$. By Schwartz’s reflection principle, $\hat{g}$ extends to a holomorphic map $\hat{g}_{*}:\hat{U}^{\prime}_{*}\to\hat{U}_{*}$ where $E_{*}$ denotes the union of a domain $E\subset\mathbb{D}^{*}$ containing $S^{1}$ in its boundary with its symmetric w.r.t. $S^{1}$ completed by $S^{1}$.
On the other hand, $\hat{g}_{*}^{-1}(S^{1})=S^{1}$, hence $\hat{g}_{*}$ has no critical points. Therefore, all pullbacks of $\hat{g}_{*}^{n}$ are well-defined in any simply connected subdomain of $(\hat{U}_{0})_{*}$. Fix such a subdomain $D$ that contains $\hat{a}$ and an arc $l$ of $S^{1}$. As $\hat{a}_{i}\to\hat{a}$ and, for each $i$, $\hat{a}_{i}$ is an attracting fixed point of the map $\hat{\psi}_{i}:\hat{Z}\to\mathbb{C}$ and $\hat{\psi}_{i}^{\prime}(\hat{a}_{i})\to 0$, we arrive at a sequence of holomorphic extensions $(\hat{\psi}_{i})_{*}$ of $\hat{\psi}_{i}$ to $D$ such that $(\hat{\psi}_{i})_{*}\to\hat{a}$ in $D$. But this cannot hold for points of the arc $l$, by the complete invariance of $S^{1}$, a contradiction
The proof for the disconnected $K$ requires new ideas and is substantially more technical. It also holds in the case of connected $K$. Let’s indicate an idea returning to the connected $K$. We don’t consider the reflection of the map $\hat{g}:\hat{U}^{\prime}\to\hat{U}$ across $S^{1}$. Instead, let us fix a simply connected domain $\Omega$ in $\mathbb{D}^{*}=\{|z|>1\}$ such that, if $\hat{Z}=\pi^{-1}(Z)$, then $\hat{a}\in\hat{Z}\subset\Omega$ and $\partial\Omega\cap S^{1}$ contains a nontrivial arc $l$ of $S^{1}$ (in fact, the set $\partial\Omega\cap S^{1}$ having a positive length would be enough). Since all branches $\hat{g}_{0}^{-n}$ are well-defined in $\Omega$, for every $i$, the local map $\hat{\psi}_{i}:\hat{Z}\to\hat{Z}$ extends to a univalent map $\hat{\psi}_{i}:\Omega\to\mathbb{C}$ where $\hat{\psi}_{i}(\hat{a}_{i})=\hat{a}_{i}$ and $\hat{\psi}_{i}^{\prime}(\hat{a}_{i})\to 0$. The point is that this implies that $\hat{\psi}_{i}$ expands the length of the arc $l$ by factors tending to $\infty$, see Appendix C. This is clearly a contradiction since all images of $l$ must lie in $S^{1}$.
@@
By Theorem A.1 and Corollary A.1, the logarithmic capacity of the completely invariant compact set $K$ is positive so Corollary B.2 applies.
Let us fix an open subset of $\mathbb{C}$ as follows:
@@
Let $\varepsilon>0$ be the distance between the set $\cup_{i=1}^{m}\partial Q_{i}$ and $K$.
For this $\varepsilon$, we find a simply connected domain $W$ as in Corollary B.2, in particular, the connected set $\pi(W)\subset\mathbb{C}\setminus K$ must lie in one and only one of the domains $Q_{1},...,Q_{m}$. W.l.o.g. one can assume that
@@
Let us joint the point $a$ and a point $q\in\partial\pi(W)\setminus K$ by a simple arc $\alpha$ which, on the one hand, lies in the range $U_{0}$ of $g$, and, on the other hand, except its end point $q$, lies outside of $\overline{\pi(W)\cup K}$ as well as outside of all others $\overline{Q_{i}}$, $2\leq i\leq m$, and also crosses the Jordan curve $\partial Q_{1}$ at a unique point. Choose a point $\hat{q}\in\partial W\cap\mathbb{D}$ such that $\pi(\hat{q})=q$ and let $\hat{\alpha}$ be the lift of the curve $\alpha$ by the covering $\pi$ that begins at $\hat{q}$. Then $\tilde{\alpha}$ is also a simple arc (in $\mathbb{D}$) with the end points $\hat{q}$ and some $\hat{a}\in\mathbb{D}$ such that $\pi(\hat{a})=a$. As $\pi$ is a local homeomorphism near $\hat{a}$, one can choose $r>0$ small enough so that there exists a topological disk $\hat{Z}$ around $\hat{a}$ such that $\pi:\hat{Z}\to Z$ where $Z=B(a,r)$ is a homeomorphism.
Note that $\hat{\alpha}\cap\overline{W}=\{\hat{q}\}$. Let us slightly "thicken" the curve $\hat{\alpha}$ to a narrow "tube" $T$ so that $\Omega:=\hat{Z}\cup T\cup W$ is a simply connected domain in $\mathbb{D}$. Then $\pi(\Omega)=Z\cup\pi(T)\cup\pi(W)$ is a domain in $\mathbb{C}\setminus K$ which is disjoint with $\overline{Q_{i}}$, for all $2\leq i\leq m$. As $\alpha$ crosses $\partial Q_{1}$ at a single point, we can also assume that $Q:=\pi(\Omega)\cup Q_{1}=\pi(T)\cup Q_{1}$ is a simply connected domain, which is disjoint with all others $Q_{2},...,Q_{m}$.
For each $i$ big enough, we can "lift" the map $\psi_{i}:Z\to Z$ to a univalent map $\hat{\psi}_{i}:\hat{Z}\to\hat{Z}$ so that $\pi\circ\hat{\psi}_{i}=\psi_{i}\circ\pi$ on $\hat{Z}$.
Claim 0. $\hat{\psi}_{i}:\hat{Z}\to\hat{Z}$ extends to a well-defined holomorphic function (also called $\hat{\psi}_{i}$) in the domain $\Omega$.
Indeed, let $\hat{\beta}$ be any curve in $\Omega$ that starts at $\hat{a}$. Then $\beta:=\pi\circ\hat{\beta}$ is a curve in $\pi(\Omega)$ that starts at $a$ and lies in $U_{0}\setminus K$. As $g^{n_{i}}:U_{-n_{i}}\setminus K\to U_{0}\setminus K$ is an unbranched covering map, the local branch $\psi_{i}:Z\to Z$ of $g^{-n_{i}}$ has an analytic continuation along the curve $\beta$, and it is lifted by the unbranched covering map $\pi:\mathbb{D}\to\hat{\mathbb{C}}\setminus K$ to an analytic continuation of $\hat{\psi}_{i}$ along the curve $\hat{\beta}$. Since the latter is any curve starting at $\hat{a}$ in the simply connected domain $\Omega$, by the Monodromy theorem, the function $\hat{\psi}_{i}$ is well-defined holomorphic continuation of $\hat{\psi}_{i}:\hat{Z}\to\hat{Z}$ to the whole domain $\Omega$.
@@
Claim 1. For each $i$, the function $\hat{\psi}_{i}:\Omega\to\mathbb{C}$ is univalent.
Indeed, assume the contrary: $\hat{\psi}_{i}(x_{1})=\hat{\psi}_{i}(x_{2})$ for some $x_{1}\neq x_{2}$ in $\Omega$. Let $\hat{l}$ be a simple curve in $\Omega$ joining $x_{1}$ and $x_{2}$. As all branches of $g^{-n_{i}}$ are inverse branches of a well-defined map $g^{n_{i}}:U_{-n}\to U_{0}$ and since $Q=\pi(\Omega)\cup Q_{1}=\pi(T)\cup Q_{1}$ is a simply connected domain, which is disjoint with some other $Q_{j}$, $j\neq 1$, the following must hold:
@@
  • an analytic continuation of some local branch of $g^{-n_{i}}$ along the closed curve $l$ maps this curve again onto some closed curve $L$,